a committee of students is made up of 6seniors and juniors. if there are 6 seniors and 8 juniors yo pick from and we need 3 from each, how many ways can the committee be made?
Found 2 solutions by oscargut, stanbon:
Answer by oscargut(2103)

(8C3)(6C3) = 56(20) = 1120
Answer: 1120

If you need more help with similar questions you can contact me now at:
mthman@gmail.com
Thanks

Answer by stanbon(75887)

a committee of students is made up of 6seniors and juniors. if there are 6 seniors and 8 juniors yo pick from and we need 3 from each, how many ways can the committee be made?
----
Ans:: 6C3 * 8C3 = 20*56 = 1020
--------------
Cheers,
Stan H.
=============